HERTA T. FREITAG

Profile

Herta Freitag was born and raised in Vienna, Austria and fell in love with mathematics at the young age of 12, despite a teacher she had for seven years and with whom she had a VERY negative relationship. There is a passage in her diary which gives the reason why she liked mathematics and wanted to become a mathematics teacher: “This is the one subject where I don’t need to memorize but can just think things out myself.” Later that year this changed to: “I don’t want to become a mathematics teacher, I want to become a good mathematics teacher.”

She received her Magister Rerum Naturalium degree, in mathematics and physics, from the University of Vienna in 1934. Later, in this country, she obtained her M.A. in 1948 and the Ph.D. in 1953, both from Columbia University.

Having come to America in 1944, she taught the first four years at Greer School in upstate New York. In 1948, she came to Hollins College, moving from instructor to full professor and chairman of the department. Jokingly, she wonders what an “empty” professor is. She “retired” in 1971, but, most fortunately, is able to continue all her mathematical activities, even sometimes (especially also for the last several years) back at Hollins in their graduate program leading to a Master of Arts in Liberal Studies degree.

For six years of her life, mathematics was taken out of her reach. As Hitler had annexed Austria in 1938, after enduring his regime for several months, she was able to await her American immigration visa in England where she “climbed up” from housemaid to governess, waitress, and – finally – even into teaching, while the bombs were raining down.

Throughout life, her early “love affair” with mathematics kept deepening. Number theory, especially recursive sequences, continued to tighten their grip on her fascination with numbers, and with problem posing and solving. She feels deeply moved when she is called a “problemist.” “I don’t call myself a mathematician,” she states, “but rather a student of mathematics, albeit quite an addicted one.” The happiest hours in her life have been spent “at the feet” of Edward Kasner and Howard Eves.

Her other interests include people, swimming and art. She regularly attends painting classes and enjoys presenting her works to her friends.